OpenStudy (anonymous):

A company has noticed its sales have started to drop exponentially. The company recorded sales in January (t = 0) as 23,000 and in March (t = 2) as 21,500. Assuming the same exponential pattern, what will its rate of decrease in sales be in June (t = 5)?

OpenStudy (jmark):

23000 = 21500e^(kt) Now, solve by: 23000/21500 = e^(k(2)) Take the natural log of both sides ln (23000/21500) = 2k Divide by 2 on both sides to obtain your k rate, which is k= -0.03372, which is your rate of decrease. Learn more about Solving Inequalities at http://math.tutorcircle.com/algebra-1/solving-inequalities.html

OpenStudy (anonymous):

@JMark that is incorrect becayse 21500 is not the original starting amount at t=0?